z is inversely proportional to x
when x=3 z=2

show that z=cy with y to the power of 9 ( where c and n are numbers and c is greater than 0(inequality)
yOU must find the values of c and n

Answer by stanbon(75887)

If z and x are inversely proportinal, z=k/x
If z=2 when x=4 then 2=k/4; then k=8, and the
equation explaining the relation is z=8/x
Your next question is almost as confusing as
your first question. Look it over and maybe
post again.
Cheers,
Stan H.